# Fusing Neural and Physical: Augment Protein Conformation Sampling with Tractable Simulations

### 16 Feb 2024 v1

@@ 3.1 Preliminaries
Molecular dynamics (MD) simulations amount to evolve the whole particle system through time $T^{\textrm{sim}}>0$  [^2] To avoid ambiguity with the time boundary $T$ in SGMs, the simulation time is denoted by $T^{\textrm{sim}}$ using superscript sim. by the physical dynamics ${d{\mathbf{x}}}={\mathbf{v}}({\mathbf{x}},\tau){d\tau}$, where the velocity field ${\mathbf{v}}\in\mathbb{R}^{N\times 3}$ can be determined by a pre-defined force field $U({\mathbf{x}})$ (aka energy or unnormalized density) via the Newtonian equation of motion: $d{\mathbf{v}}({\mathbf{x}},\tau)=\bar{{\mathbf{m}}}\nabla_{\mathbf{x}}U({\mathbf{x}})~{}d\tau$, where $\bar{{\mathbf{m}}}$ is the inverse mass of atoms in the system. At each time step, the position and velocity may be updated by a symplectic integrator such as Verlet (Verlet, 1967), yielding a simulation trajectory.
@@ 3.2 Neighborhood exploitation with short MD
Figure 2: Illustration of the conformation sampling of Str2Str (Lu et al., 2024). In (a), sampling is performed in two steps to obtain i.i.d. samples of the target protein, where the energy landscape (or information of the force field) is unknown and colored gray. In (b), a hypothetical zoom-in neighborhood of a sample is shown. Due to the generally complicated landscape of force field, the samples directly generated by Str2Str are probably not energy-optimal. Short MD simulation can be run to obtain locally equilibrium samples and can be used to fine-tune the pre-trained sampler.
@@ 4 Experiments
To validate the effectiveness of the proposed fine-tuning pipeline, we evaluate the sampling performance on the fast-folding protein set (Lindorff-Larsen et al., 2011). For baselines, we compare the fine-tuned Str2Str (namely Str2Str-FT) against: AF2-RAVE (Vani et al., 2023), EigenFold (Jing et al., 2023), idpGAN (Janson et al., 2023), and Str2Str (pre-trained only) (Lu et al., 2024). The size of sampled ensembles is aligned to be 1,000 to make fair comparison. We follow the evaluation metrics used in Lu et al. (2024) and list them as below [^3] We omit the computationally extensive diversity metrics which involve pairwise TM-score since the distribution distance metrics (JS) can also reflect diversity mismatch.: (1) Validity: Val-Clash, Val-Bond evaluates whether the conformations obey basic geometric constraints w.r.t. steric clash and bonding association. These metrics are the higher the better ($\uparrow$). (2) Distance: JS-PwD, JS-TIC, JS-Rg assess the distance between the model distribution and reference distribution, which includes the pairwise distance (PwD), time-lagged independent components (TIC) (Naritomi & Fuchigami, 2011; Pérez-Hernández et al., 2013) and radius of gyration (Rg), which are the lower the better ($\downarrow$).
@@
In this study, we limit the few-shot budget of MD production simulations to be 100ns in total, which is computationally tractable and lasts only several GPU hours. Under this budget, we firstly sample for each target 100 conformations from a pre-trained Str2Str followed by 100 short simulations per 1 ns long. As shown in Table 1, both the sampler with neighborhood exploitation (Str2Str-NE) and the fine-tuned one (Str2Str-FT) achieve the state-of-the-art performance across all three distance metrics while keeping good validity.
@@ 5 Conclusion
In this work, we explore the few-shot setting of protein conformation sampling by combining score-based generative models and physical MD simulations. Generative models explore the conformation space efficiently while the physical force field refine the unreasonable samples, achieving good exploration-exploitation balance. The fine-tuned diffusion sampler has shown significant improvement on conformation sampling, which exemplified the domain adaption of Str2Str with efficient short simulations. Future works can focus on the study of more effective fine-tuning strategies and enabling the sampling approximately from Boltzmann distribution.
@@ A.2 Fine-tuning dataset
To acquire the fine-tuning dataset (ES) for each target, we firstly generate the seeding structure using ESMFold (Lin et al., 2023) with default configuration and ran Str2Str as described in Lu et al. (2024) to generate 100 conformations from the seeding structure. Then MD simulations was performed in parallel initialized respectively from these conformations according to the protocol above in Appendix A.1. We ignored those simulation threads failed due to possible bad initial geometries. The resulting conformations as well as their associated potential energies were saved to construct the fine-tuning dataset. We split the dataset into training and validation set with ratio $0.95:0.05$.
@@ A.3 Setup of training and inference
To optimize the network parameters of the pre-trained Str2Str, the Adam optimizer (Kingma & Ba, 2014) was used with $\rm{lr}=10^{-4}$ and $\beta_{1}=0.9,\beta_{2}=0.999$. As scheduler, the learning rate was reduced in the factor of 0.1 when the loss has stopped decreasing for 10 epochs. The training process was set to be at least 100 epochs and mediated the early stopping strategy by monitoring the validation loss with 10 epochs patience. The fine-tuning time for different targets was on-average $\sim$ 2 GPU hours on the NVIDIA Tesla V100-SXM2-32GB GPU. For simplicity, we use uniform weighting in this study for the training objective, i.e. $\omega(x,t)\equiv 1$. Other model configurations were kept the same as in Lu et al. (2024). After fine-tuning, we sampled the target protein from the prior distribution $p_{T}({\mathbf{x}})$ with discretization of 1,000 timesteps by reversing the diffusion process via Langevin dynamics. In practice, a minimum time $t_{\epsilon}=0.01$ instead of zero is used for numerical stability. The SDE mode of sampling was used for both Str2Str and Str2Str-FT.
A.4 Fast-folding Targets.
@@
A.5 Extended experimental results
@@
To illustrate the efficiency of proposed few-shot sampling over conventional MD simulation, we here profile the wall clock running time of each type of ”sampler” for the benchmark targets, shown in Table 3. Note that the runtime can vary from different proteins due to the different number of atoms. The profiling results are all reported as wall clock hours based on 4 NVIDIA Tesla V100-SXM2-32GB GPUs.
Figure 3: The runtime profile of different samplers used in this study

### 11 Mar 2024 v2

@@ 3.1 Preliminaries
Molecular dynamics (MD) simulations amount to evolve the whole particle system through time $T^{\textrm{sim}}>0$ by the physical dynamics ${d{\mathbf{x}}}={\mathbf{v}}({\mathbf{x}},\tau){d\tau}$, where ${\mathbf{v}}\in\mathbb{R}^{N\times 3}$ is the velocity field. With some force field defining $U({\mathbf{x}})$ as potential energy, the velocity can be updated through some dynamics such as Newtonian equation [^2] Sometimes referred to as Hamiltonian dynamics. Besides, stochastic dynamics can also be applied such as Langevin dynamics to update the velocity states.: $d{\mathbf{v}}({\mathbf{x}},\tau)=\bar{{\mathbf{m}}}\nabla_{\mathbf{x}}U({\mathbf{x}})~{}d\tau$, where $\bar{{\mathbf{m}}}$ is the inverse mass of atoms in the system. At each time step, the position and velocity can be updated by an integrator such as Verlet (Verlet, 1967), yielding a simulation trajectory.
@@ 3.2 Neighborhood exploitation with short MD
Figure 2: Illustration of the conformation sampling scenario of Str2Str. In (a), sampling is performed in two steps to obtain independent samples of the target protein, where the energy landscape (or information of the force field) is unknown and colored in gray. In (b), a hypothetical zoom-in neighborhood of a sample is shown. Due to the complex conformation landscape, the samples directly generated by Str2Str are probably not potential energy-optimal. Short MD simulation can be run to obtain locally equilibrated samples which can be used to fine-tune the pre-trained sampler.
@@ 4 Experiments
To validate the effectiveness of the proposed fine-tuning pipeline, we evaluate the sampling performance on the fast-folding protein set (Lindorff-Larsen et al., 2011). For baselines, we compare the fine-tuned Str2Str (namely Str2Str-FT) against: MSA subsampling (Del Alamo et al., 2022), EigenFold (Jing et al., 2023), idpGAN (Janson et al., 2023), and Str2Str (pre-trained only) (Lu et al., 2024). The size of sampled ensembles is aligned to be 1,000 to make fair comparison. We follow the evaluation metrics used in Lu et al. (2024) and list them as below: (1) Validity: Val-Clash, Val-Bond evaluates whether the conformations obey basic geometric constraints w.r.t. steric clash and bonding association. These metrics are the higher the better ($\uparrow$). (2) Distance: JS-PwD, JS-TIC, JS-Rg assess the distance between the model distribution and reference distribution, which includes the pairwise distance (PwD), time-lagged independent components (TIC) (Naritomi & Fuchigami, 2011; Pérez-Hernández et al., 2013) and radius of gyration (Rg), which are the lower the better ($\downarrow$).
@@
In this study, we limit the few-shot budget of MD production simulations to be 100ns in total, which is computationally tractable and lasts only several GPU hours. Under this budget, we firstly sample for each target 100 conformations from a pre-trained Str2Str followed by 100 short simulations per 1 ns long. As shown in Table 1, both the sampler with neighborhood exploitation (Str2Str-NE) and the fine-tuned version (Str2Str-FT) achieve the state-of-the-art performance across all three distance metrics while keeping good validity.
@@ 5 Conclusion
In this work, we explore the few-shot settings of protein conformation sampling by combining score-based generative models and physical MD simulations. Generative models explore the conformation space efficiently while the physical force field refine the unreasonable samples, achieving good exploration-exploitation balance. The fine-tuned diffusion sampler has shown significant improvement on conformation sampling, which exemplified the domain adaption of Str2Str with efficient short simulations. Future works can focus on the study of more effective fine-tuning strategies to enable sampling approximately from Boltzmann distribution.
Acknowledgments
This project is supported by Twitter, Intel, the Natural Sciences and Engineering Research Council (NSERC) Discovery Grant, the Canada CIFAR AI Chair Program, Samsung Electronics Co., Ltd., Amazon Faculty Research Award, Tencent AI Lab Rhino-Bird Gift Fund, an NRC Collaborative R&D Project (AI4D-CORE-06) as well as the IVADO Fundamental Research Project grant PRF-2019-3583139727.
@@ A.2 Fine-tuning dataset
To acquire the fine-tuning dataset (ES) for each target, we firstly generate the seeding structure using ESMFold (Lin et al., 2023) with default configuration and ran Str2Str as described in Lu et al. (2024) to generate 100 conformations from the seeding structure. Then MD simulations was performed in parallel initialized respectively from these conformations according to the protocol above in Appendix A.1. We ignored those simulation threads failed due to possible bad initial geometries. The resulting conformations as well as their associated potential energies were saved to construct the fine-tuning dataset. We randomly split the dataset into training and validation set with ratio $0.95:0.05$ while the latter is used for early stopping of training.
@@ A.3 Setup of training and inference
To optimize the network parameters of the pre-trained Str2Str, the Adam optimizer (Kingma & Ba, 2014) was used with $\rm{lr}=10^{-4}$ and $\beta_{1}=0.9,\beta_{2}=0.999$. As scheduler, the learning rate was reduced in the factor of 0.1 when the loss has stopped decreasing for 10 epochs. The training process was set to be at least 100 epochs and mediated the early stopping strategy by monitoring the validation loss with 10 epochs patience. The fine-tuning time for different targets was on-average $\sim$ 2 GPU hours on the NVIDIA Tesla V100-SXM2-32GB GPU. For simplicity, we use uniform weighting in this study for the training objective, i.e. $\omega(x,t)\equiv 1$. Other model configurations and evaluation pipelines were kept the same as in Lu et al. (2024). After fine-tuning, we sampled the target protein from the prior distribution $p_{T}({\mathbf{x}})$ with discretization of 1,000 timesteps by reversing the diffusion process via Langevin dynamics. In practice, a minimum time $t_{\epsilon}=0.01$ instead of zero is used for numerical stability. The SDE mode of sampling was used for both Str2Str and Str2Str-FT.
A.4 Fast-folding Targets
@@
Appendix B Evaluation metrics
[…]
Appendix C Extended experimental results
@@
To illustrate the efficiency of proposed few-shot sampling over conventional MD simulation, we here profile the wall clock running time of each type of ”sampler” for each benchmark target, shown in Table 3. Note that the runtime can vary from different proteins due to the different number of atoms. The profiling results are all reported as the wall clock hours based on 4$\times$ NVIDIA Tesla V100-SXM2-32GB GPUs.
Figure 3: The runtime profile of different samplers across each fast folding target used in this study.
Appendix D Limitations and Discussion
Similar to Str2Str (Lu et al., 2024), the proposed methods cannot guarantee the sampled ensemble are Boltzmann distributed (Noé et al., 2019) due to the nature of local simulation. It is not designed for replacing canonical MD simulations for the study of protein dynamics. Incorporating short MD simulations in the Str2Str sampler takes the advantages from both physical and neural sides: the neural sampler (eg., diffusion) is good at proposing data-like conformation hypothesis very efficiently in the vast conformation space while the physical simulation can equilibrate the samples onto some local minima. Both neighborhood exploitation (NE) and fine-tuning (FT) provide a simple yet effective way to refine the resulting conformations sampled from a neural sampler.
From another perspective, neighborhood exploitation coincides with the purpose of enhanced sampling methods. Enhanced sampling methods aim to overcome the energy barrier and accelerate the exploration of long MD simulations. For example, the umbrella sampling (Torrie & Valleau, 1977), and parallel tempering (synonymous replica exchange molecular dynamics, REMD) (Hansmann, 1997; Sugita & Okamoto, 1999; Swendsen & Wang, 1986). The neural sampler such as Str2Str can provide diverse conformation seeds to augment the exploration of MD simulation similar to Vani et al. (2023). We find it a promising research direction to develop the neural enhanced sampling methods in the future.